# XLand-100B: A Large-Scale Multi-Task Dataset for In-Context Reinforcement Learning

### v1

@@ 6 Limitations and Future Work
Acknowledgments and Disclosure of Funding
At the time of writing, all authors were employed by AIRI or/and Tinkoff. All computational resources were provided by AIRI and Tinkoff.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 4.
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A] We provide reasons for this answer in Appendix B
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes] We have read the ethics guidelines and checked our paper satisfies them
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 4.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] Refer to Appendix I and LABEL:apndx:hyper.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] They are showed and explained in the captions.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] Refer to Appendix I and Appendix J.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A]
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A License
We distribute our dataset under CC BY-SA 4.0 license.
Appendix B General Ethic Conduct and Potential Negative Societal Impact
To the best of our knowledge, our work does not present any direct potential negative societal impact, besides the carbon emissions produced by the GPU accelerators during datasets collection.
Appendix C Downloading the Datasets
@@
Both XLand-100B and XLand-Trivial-20B datasets hosted on public S3 bucket and freely available for everyone under CC BY-SA 4.0 Licence. We advise starting with Trivial dataset for debugging due to smaller size and faster downloading time.
Datasets can be downloaded with the curl (or any other similar) utility: curl -L -o xland-trivial-20b.hdf5 https://sc.link/A4rEW for smaller version and curl -L -o xland-100b.hdf5 https://sc.link/MoCvZ for main version.
Appendix D What is Inside Dataset?
@@
Appendix E Hyperparameters
@@
Appendix F Compression Chunk Size Tuning
Compression Chunk size Throughput
None None 1,549,619
gzip None 173,895
gzip 256 423,513
gzip 512 549,397
gzip 1024 666,152
gzip 2048 768,706
gzip 4096 749,851
gzip 8192 737,646
Appendix G Additional Figures of Data Collection
Figure 10: Final return by number of rules on 256 tasks for policies trainef from scratch or fine-tuned from multi-taks pre-trained checkpoints.
Appendix H Additional Figures of Data Evaluation
Appendix I AD Implementation
Table 9: Time for training and evaluation.
sequence length 1024 2048 4096
train 15 hrs 10 hrs 11.5 hrs
eval 20 min 30 min 50 min
Appendix J DPT Implementation
We also show DPT training logs. Trivial: wandb; medium: wandb
J.1 Dark Key-To-Door Realization
Table 10: Q-Learning Hyperparameters
Hyperparameter Value
Num. Train Goals 2424
Num. Histories 5000
Num. Updates 50,000
Learning Rate 3e-4
Appendix K Additional Figures of AD Performance
Appendix L Additional Figures of DPT Performance

### v2

@@ 6 Limitations and Future Work
Ethics Statement
Legal Compliance and Licensing. The dataset will be released under the CC BY-SA 4.0 license, and the accompanying code will be available under the Apache 2.0 license. We have ensured that our data and code releases comply with all relevant legal and ethical guidelines.
Data Privacy and Security. The dataset does not include any human subjects or personally identifiable information. All data are generated by artificial agents interacting within a simulated environment. Consequently, there are no privacy or security concerns associated with the dataset.
Potential Misuse and Harmful Applications. The dataset and methods are intended to advance research in in-context reinforcement learning. We do not anticipate any immediate risks of misuse or harmful applications arising from our work. Nonetheless, we encourage users to apply the dataset responsibly and adhere to ethical standards in their research.
Environmental Impact. Collecting the dataset required substantial computational resources, totaling approximately 50,000 GPU hours. We acknowledge the environmental footprint associated with large-scale computations. By openly releasing this dataset, we aim to minimize redundant data collection efforts by other researchers, potentially reducing the overall environmental impact in this field.
Reproducibility Statement
To ensure reproducibility we share the important details throughout the paper and appendix. We explain how we pretrain, finetune models for collecting data and the post-processing step in Section 4.2, as well as discussing the dataset details (size, compression ratios, etc.) in Appendix B and Appendix C. For experiments section, we discuss the details of baselines implementation in Appendix F and Appendix G. We provide hyperparameters for collecting the dataset and for baselines in Appendix O. We also release the codebase with tools for creating and expanding the dataset in the following repository: xland-minigrid-datasets.
Appendix A Downloading the Datasets
@@
Both XLand-100B and XLand-Trivial-20B datasets hosted on public S3 bucket and are freely available for everyone under CC BY-SA 4.0 Licence. We advise starting with Trivial dataset for debugging due to smaller size and faster downloading time. Datasets can be downloaded with the curl (or any other similar) utility.
# XLand-Trivial-20B, approx 60GB size
curl -L -o xland-trivial-20b.hdf5 https://tinyurl.com/trivial-10k
# XLand-100B, approx 325GB size
curl -L -o xland-100b.hdf5 https://tinyurl.com/medium-30k
Appendix B What is Inside Dataset?
@@
Appendix C Compression Chunk Size Tuning
Compression Chunk size Throughput
None None 1,549,619
gzip None 173,895
gzip 256 423,513
gzip 512 549,397
gzip 1024 666,152
gzip 2048 768,706
gzip 4096 749,851
gzip 8192 737,646
Appendix D Additional Figures of Data Collection
Figure 10: Final return by number of rules on 256 tasks for policies trained from scratch or fine-tuned from multi-taks pre-trained checkpoints.
Appendix E Additional Figures of Data Evaluation
Appendix F AD Implementation
Table 5: Time for training and evaluation.
sequence length 1024 2048 4096
train 15 hrs 10 hrs 11.5 hrs
eval 20 min 30 min 50 min
Appendix G DPT Implementation
We also show DPT training logs. Trivial: wandb; medium: wandb
Appendix H On DPT Limitations in POMDP
Table 6: Q-Learning Hyperparameters
Hyperparameter Value
Num. Train Goals 2424
Num. Histories 5000
Num. Updates 50,000
Learning Rate 3e-4
Appendix I Details of In-Context Evaluation
The evaluation process resembles the standard in-weight learning, with the key difference being that the learning itself happens during evaluation. When interacting with the environment, the agent populates the Transformer’s context with the latest observations, actions and rewards. At the start of the evaluation, the context is empty. Note that the agent’s context is cross-episodic, which makes it possible for the agent to access transitions in previous episodes. We report the cumulative reward that the agent achieved at the end of each episode. We run evaluation for each model for 500 episodes, reporting mean return across 1024 unseen tasks with standard deviation across 3 seeds. We claim that the in-context emerged when the mean return rises up until some level. This means the agent improves its policy from episode to episode, learning how to solve a task.
Appendix J Additional Figures of AD Performance
Appendix K Additional Figures of DPT Performance
Appendix L Rulesets Vizualization
Figure 21: Rulesets with increasing difficulty (i.e. with 3, 5, and 9 rules respectively) sampled from medium-1m benchmark. They can be accessed by ID’s via the XLand-MiniGrid package.
Appendix M Grid Layouts Used for Dataset Collection
Figure 22: Layout used for XLand-Trivial-20B. Corresponds to XLand-MiniGrid-R1-9x9 environment from the XLand-MiniGrid suite. Game objects are task dependent.
Figure 23: Layout used for XLand-100B. Corresponds to XLand-MiniGrid-R1-13x13 environment from the XLand-MiniGrid suite. Game objects are task dependent.
Appendix N Additional experiments on number of tasks
We run additional experiments with different number of tasks to show how it can affect the performance. Experiments were done with the sequence length of 2048. To make the number of gradient updates closer to the original experiments, we progressively increased the number of epochs for each smaller subset. Note that the return increases from 100 goals to 3k goals, indicating the importance of scaling the number of goals in the dataset. Running the AD for a single epoch on full goals dataset is clearly not enough, however, a large scale training is too computationally expensive and it is left for future research.
Figure 24: Training AD for different number of goals.
Appendix O Hyperparameters